### v1

# A Wasserstein perspective of Vanilla GANs

A Wasserstein perspective of Vanilla GANs
Lea Kunkel and Mathias Trabs
(Karlsruhe Insitute of Technology)
The empirical success of Generative Adversarial Networks (GANs) caused an increasing interest in theoretical research. The statistical literature is mainly focused on Wasserstein GANs and generalizations thereof, which especially allow for good dimension reduction properties. Statistical results for Vanilla GANs, the original optimization problem, are still rather limited and require assumptions such as smooth activation functions and equal dimensions of the latent space and the ambient space. To bridge this gap, we draw a connection from Vanilla GANs to the Wasserstein distance. By doing so, existing results for Wasserstein GANs can be extended to Vanilla GANs. In particular, we obtain an oracle inequality for Vanilla GANs in Wasserstein distance. The assumptions of this oracle inequality are designed to be satisfied by network architectures commonly used in practice, such as feedforward ReLU networks. By providing a quantitative result for the approximation of a Lipschitz function by a feedforward ReLU network with bounded Hölder norm, we conclude a rate of convergence for Vanilla GANs as well as Wasserstein GANs as estimators of the unknown probability distribution.
Keywords: Generative adversarial networks, rate of convergence, oracle inequality, Wasserstein distance, distribution estimation
MSC 2020: 62E17, 62G05, 68T07
1 Introduction
Generative Adversarial Networks (GANs) have attracted much attention since their introduction by Goodfellow et al., (2014), initially due to impressive results in the creation of photorealistic images. Meanwhile, the areas of application have expanded far beyond this, and GANs serve as a prototypical example of the rapidly evolving research area of generative models.
The Vanilla GAN as constructed by Goodfellow et al., (2014) relies on the minimax game
$\inf_{G\in\mathcal{G}}\sup_{D\in\mathcal{D}}\mathbb{E}\big{[}\log D(X)+\log\big{(}1-D(G(Z))\big{)}\big{]},$ (1)
to learn an unknown distribution $\mathbb{P}^{*}$ of the random variable $X$. The generator $G$ chosen from a set $\mathcal{G}$ applied to the latent random variable $Z$ aims to mimic the distribution of $X$ as closely as possible. The discriminator $D,$ chosen from a set $\mathcal{D},$ has to distinguish between real and fake samples.
The optimization problem (1) is motivated by the Jensen-Shannon divergence. Generalizations of the underlying distance have led to various extensions of the original GAN, such as $f$-GANs (Nowozin et al., (2016)). More famously, Wasserstein GANs (Arjovsky et al., (2017)), characterized by
$\inf_{G\in\mathcal{G}}\sup_{W\in\operatorname*{Lip}(1)}\mathbb{E}\big{[}W(X)-W(G(Z))\big{]},$ (2)
are obtained by replacing the Jensen-Shannon divergence by the Kantorovich dual of the Wasserstein distance. Here, $\operatorname*{Lip}(1)$ denotes the set of all Lipschitz continuous functions with Lipschitz constant bounded by one. This approach can be generalized to Integral Probability Metrics (IPM, Mueller, (1997)) GANs, see Liang, (2021).
The analysis of Wasserstein GANs can exploit the existing theory on the Wasserstein distance. The latter has a long record of research, particularly in the context of optimal transport (Villani,, 2008) but also in machine learning, see Torres et al., (2021) for an overview. In contrast, Vanilla GANs and the Jensen-Shannon divergence have been studied less extensively, and fundamental questions have not been settled. In particular, all statistical results for Vanilla GANs require the same dimension of the latent space and the target space which is in stark contrast to common practice. Another algorithmic drawback of Vanilla GANs highlighted by Arjovsky & Bottou, (2017) is that an arbitrarily large discriminator class prevents the generator from learning. The Jensen-Shannon divergence between singular measures is by definition maximal. Therefore, we cannot expect convergence in a dimension reduction setting. In practice, however, Vanilla GANs have worked in various settings. Thus, using neural networks as a discriminator class must be advantageous compared to the set of all measurable functions. This empirical fact is supported by the numerical results by Farnia & Tse, (2018) who impose a Lipschitz constraint on the discriminator class.
Our contribution.
Our work aims to bridge the gap between Vanilla GANs and Wasserstein GANs while addressing the theoretical limitations of the former ones. By imposing a Lipschitz condition on the discriminator class in (1), we recover a Wasserstein-like behavior. As a main result, we can derive an oracle inequality for the Wasserstein distance between the true data generating distribution and its Vanilla GAN estimate. In particular, this allows us to transfer key features, such as dimension reduction, known from the statistical analysis of Wasserstein GANs. We show that the statistical error of the modified Vanilla GAN depends only on the dimension of the latent space, independent of the potentially much larger dimension of the feature space $\mathcal{X}$. Thus Vanilla GANs can avoid the curse of dimensionality. Such properties are well known from practice, but cannot be verified by the classical Jensen-Shannon analysis. On the other hand the derived rate of convergence for the Vanilla GAN is slower than for Wasserstein GANs which is in line with the empirical advantage of Wasserstein GANs.
We then consider the most relevant case where the classes $\mathcal{G}$ and $\mathcal{D}$ are parameterized by neural networks. Using our previous results, we derive an oracle inequality that depends on the network approximation errors for the best possible generator and the optimal Lipschitz discriminator. To bound the approximation error of the discriminator, we replace the Lipschitz constraint on the networks with a less restrictive Hölder constraint. Building on Gühring et al., (2020), we prove a novel quantitative approximation theorem for Lipschitz functions using ReLU neural networks with bounded Hölder norm. As a result we obtain the rate of convergence $n^{-\alpha/2d^{*}},\alpha\in(0,1),$ with latent space dimension $d^{*}\geq 2$ for sufficiently large classes of networks. Additionally, our approximation theorem allows for an explicit bound on the discriminator approximation error for Wasserstein-type GANs, which achieve the rate $n^{-\alpha/d^{*}},\alpha\in(0,1)$.
Related work.
The existence and uniqueness of the optimal generator for Vanilla GANs is shown by Biau et al., (2018) under the condition that the class $\mathcal{G}$ is convex and compact. They also study the asymptotic properties of Vanilla GANs. Belomestny et al., (2021) have shown a non-asymptotic rate of convergence in the Jensen-Shannon divergence for Vanilla GANs with neural networks under the assumption that the density of $\mathbb{P}^{*}$ exists and that the generator functions are continuously differentiable.
In practice, however, the Rectifier Linear Unit activation function (ReLU activation function) is commonly used (Aggarwal,, 2018, p.13). The resulting neural network generates continuous piecewise linear functions. Therefore, the convergence rate of Belomestny et al., (2021) combined with Belomestny et al., (2023) is not applicable to this class of functions.
The statistical analysis of Wasserstein GANs is much better understood. Biau et al., (2021) have studied optimization and asymptotic properties. Liang, (2021) has shown error decompositions with respect to the Kullback-Leibler divergence, the Hellinger distance and the Wasserstein distance. The case where the unknown distribution lies on a low-dimensional manifold is considered in Schreuder et al., (2021) as well as Tang & Yang, (2022). The latter also derive minmax rates in a more general setting using the Hölder metric. Assuming that the density function of $\mathbb{P}^{*}$ exists, Liang, (2017) has shown a rate of convergence in Wasserstein distance with ReLU activation function with a factor growing exponentially in the depth of the network. Theoretical results including sampling the latent distribution in addition to dimension reduction have been derived by Huang et al., (2022), who have also shown a rate of convergence in a slightly more general setting (using the Hölder metric) using ReLU networks whose Lipschitz constant grows exponentially in the depth. A rate of convergence using the total variation metric and leaky ReLU networks has been shown in Liang, (2021).
Convergence rates with respect to the Wasserstein distance have been studied by Chae, (2022) and Chen et al., (2020). Up to a logarithmic factor, optimal rates in the Hölder metric were obtained by Stéphanovitch et al., (2023) using smooth networks. In a similar setting, Chakraborty & Bartlett, (2024) discussed several methods for dimension reduction. Recently, Suh & Cheng, (2024) have reviewed the theoretical advances in Wasserstein GANs.
Ensuring Lipschitz continuity of the discriminator class is the essential property of Wasserstein GANs. Lipschitz-constrained neural networks and their empirical success are subject of ongoing research
(Khromov & Singh, (2023), in context of GANs see Than & Vu, (2021)). Implementations of Wasserstein GANs have evolved from weight clipping (Arjovsky et al.,, 2017) to penalizing the objective function (Gulrajani et al.,, 2017; Wei et al.,, 2018; Zhou et al.,, 2019; Petzka et al.,, 2017; Miyato et al.,, 2018; Asokan & Seelamantula,, 2023), which heuristically leads to networks with bounded Lipschitz constants. Farnia & Tse, (2018) use an objective function that combines Wasserstein and Vanilla GANs.
Outline.
In Section 2 we introduce the Vanilla GAN distance, which characterizes the optimization problem (1). In Section 3, we investigate the relation between the Vanilla GAN distance and the Wasserstein distance. We show that the distances are compatible to each other while not being equivalent. Using this relation, we derive an oracle inequality for the Vanilla GAN in Section 4, where $\mathcal{G}$ is an arbitrary set and $\mathcal{D}$ is a set of Lipschitz functions. We show that Vanilla GANs can avoid the curse of dimensionality. In Section 5 we consider the situation where $\mathcal{G}$ and $\mathcal{D}$ consist of neural networks. Here we relax the Lipschitz condition to a $\alpha$-Hölder condition and prove a quantitative result for the approximation of a Lipschitz function by a feedforward ReLU network with bounded Hölder norm. We then prove a convergence rate for the Vanilla GAN with network generator and discriminator. All proofs are deferred to section 7.
2 The Vanilla GAN distance
Let us first fix some notation. We equip $\mathbb{R}^{d}$ with the $\ell_{p}$-norm $|x|_{p}$, $1\leq p\leq\infty$, and denote the number of nonzero entries of a $k\times l$ matrix $A,$ where $k,l\in\mathbb{N},$ by $|A|_{\ell^{0}}\coloneqq|\{(i,j)\colon A_{ij}\neq 0\}|.$ For ease of notation we abbreviate for $x\in(0,\infty)$
$[x]^{1;1/2}\coloneqq\max\{x,\sqrt{x}\}.$ (3)
For $\Omega\subset\mathbb{R}^{d}$ and $f\colon\Omega\rightarrow\mathbb{R}$ we define $\|f\|_{\infty,\Omega}:=\operatorname{ess\ sup}\{|f(x)|:x\in\Omega\}.$ We denote the set of bounded Lipschitz functions by
$\operatorname*{Lip}(L,B,\Omega)\coloneqq\{f\colon\Omega\rightarrow\mathbb{R}\mid\|f\|_{\infty,\Omega}\leq B,\ |f(x)-f(y)|\leq L|x-y|_{p},\ \forall x,y\in\Omega\}.$
The set of unbounded Lipschitz functions is denoted by $\operatorname*{Lip}(L,\Omega)\coloneqq\operatorname*{Lip}(L,\infty,\Omega).$ By Rademacher’s theorem, a Lipschitz function is differentiable almost everywhere. For $\alpha\in(0,1]$ we define the $\alpha$-Hölder norm by
$\|f\|_{\mathcal{H}^{\alpha}(\Omega)}:=\max\left\{\|f\|_{{\infty}},\;\underset{x,y\in\Omega}{\operatorname{ess\;sup}}\;\frac{|f(x)-f(y)|}{|x-y|_{p}^{\alpha}}\right\}$ (4)
and the $\alpha$-Hölder ball of functions with Hölder constant less or equal than $\Gamma>0$ as
$\mathcal{H}^{\alpha}(\Gamma,\Omega):=\Big{\{}f\colon\Omega\to\mathbb{R}\leavevmode\nobreak\ \Big{|}\leavevmode\nobreak\ \|f\|_{\mathcal{H}^{\alpha}(\Omega)}<\Gamma\Big{\}}.$ (5)
In particular, $\operatorname*{Lip}(L,B,\Omega)\subseteq\mathcal{H}^{\alpha}(\max(L,2B),\Omega)$ for any $\alpha\in(0,1).$ We omit the domain $\Omega$ in our notation if $\Omega=(0,1)^{d}$.
We observe an i.i.d. sample $X_{1},...,X_{n}\sim\mathbb{P}^{*}$ with values in $\mathcal{X}\coloneqq(0,1)^{d}$. On another space $\mathcal{Z}\coloneqq(0,1)^{d^{*}}$, called the latent space, we choose a latent distribution $\mathbb{U}$. Unless otherwise specified, $X\sim\mathbb{P}^{*}$ and $Z\sim\mathbb{U}$. Throughout, the generator class $\mathcal{G}$ is a compact nonempty set of measurable functions from $\mathcal{Z}$ to $\mathcal{X}$. For $G\in\mathcal{G}$ the distribution of the random variable $G(Z)$ is denoted by $\mathbb{P}^{G(Z)}.$
Typically the discriminator class consists of functions mapping to $\mathbb{R}$ concatenated to a sigmoid function that maps into $(0,1)$ to account for the classification task. This is especially the case for standard classification networks. The most common sigmoid function used for this purpose is the logistic function $x\mapsto(1+e^{-x})^{-1},$ which we fix throughout. Together with a shift by $\log 4$, we can rewrite the Vanilla GAN optimization problem (1) as
$\inf_{G\in\mathcal{G}}\mathsf{V}_{\mathcal{W}}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})$ (6)
in terms of the Vanilla GAN distance between probability measures $\mathbb{P}$ and $\mathbb{Q}$
$\mathsf{V}_{\mathcal{W}}(\mathbb{P},\mathbb{Q})\coloneqq\sup_{W\in\mathcal{W}}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}\Big{[}-\log\Big{(}\frac{1+e^{-W(X)}}{2}\Big{)}-\log\Big{(}\frac{1+e^{W(Y)}}{2}\Big{)}\Big{]},$ (9)
where $\mathcal{W}$ is a set of measurable functions $W\colon\mathcal{X}\to\mathbb{R}$. As long as $0\in\mathcal{W},$ we have that $\mathsf{V}_{\mathcal{W}}\geq 0.$
To choose the generator $\hat{G}_{n}$ as the empirical risk minimizer, the unknown distribution $\mathbb{P}^{*}$ in (6) must be replaced by the empirical distribution $\mathbb{P}_{n}$ based on the observations $X_{1},...,X_{n}$. In practice, the expectation with respect to $Z\sim\mathbb{U}$ is replaced by an empirical mean too, which we omit for the sake of simplicity. Following Huang et al., (2022), the next and all subsequent results easily extend to the corresponding setting.
The following error bound in terms of the Vanilla GAN distance provides an error decomposition for the empirical risk minimizer of the Vanilla GAN.
Lemma 2.1.
Assume that $\mathcal{G}$ is chosen such that a minimum exists. Let $\mathcal{W}$ be symmetric, that is, for $W\in\mathcal{W},-W\in\mathcal{W}.$ For
$\hat{G}_{n}\in\underset{G\in\mathcal{G}}{\arg\min}\;\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})$ (10)
we have that
$\mathsf{V}_{\mathcal{W}}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})\leq$ $\min_{G\in\mathcal{G}}\mathsf{V}_{\mathcal{W}}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})+2\sup_{W\in\operatorname*{Lip}(1)\circ\mathcal{W}}\frac{1}{n}\sum_{i=1}^{n}\big{(}W(X_{i})-\mathbb{E}[W(X)]\big{)}.$ (11)
The first term in (11) is the error due to the approximation capabilities of the class $\mathcal{G}.$ The second term refers to the stochastic error due to the amount of training data. As $\mathcal{W}$ is symmetric, the stochastic error is non-negative. Both error terms depend on the discriminator class $\mathcal{W}$. Large discriminator classes lead to finer discrimination between different probability distributions and thus to a larger approximation error term. Similarly, the stochastic error term will increase with the size of $\mathcal{W}$. The cost of small classes $\mathcal{W}$ is a less informative loss function on the left side of (11).
If $\mathcal{W}$ is the set of all measurable functions, the analysis by Goodfellow et al., (2014, Theorem 1) shows that the Vanilla GAN distance is equivalent to the Jensen-Shannon distance shifted by $\log(4)$. Arjovsky & Bottou, (2017) have elaborated on the theoretical and practical disadvantages of this case. Similar to the total variation distance or the Hellinger distance, the Jensen-Shannon divergence is not compatible with high-dimensional settings because it cannot distinguish between different singular measures. Therefore, we need a weaker distance and thus restrict $\mathcal{W}$.
The key insight of Wasserstein GANs (2) is that this particular drawback of the Jensen-Shannon distance is can be solved by the Wasserstein distance. The latter is a metric on the space of probability distributions with finite first moment and meterizes weak convergence in this space (Villani,, 2008, Theorem 6.9). Let $\mathbb{P}$ and $\mathbb{Q}$ be two probability distributions on the same measurable space $(\Omega,\mathcal{A})$, the Wasserstein-$1$ distance is defined as
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})\coloneqq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1)\\ W(0)=0\end{subarray}}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Y)]=\sup_{W\in\operatorname*{Lip}(1)}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Y)].$ (18)
Bounds for weaker metrics, such as the Kolmogorov or Levy metric, can be easily derived from the bounds in the Wasserstein metric under weak conditions, see e.g. Gibbs & Su, (2002).
Therefore, we choose $\mathcal{W}=\operatorname*{Lip}(L)$ for some $L\geq 1$ in Lemma 2.1. In this case the following result shows that the existence of an empirical risk minimizer is guaranteed as soon as $\mathcal{G}$ is compact.
Lemma 2.2.
Assume $\mathcal{G}$ is compact with respect to the supremum norm. The map $T\colon\mathcal{G}\rightarrow\mathbb{R}_{\geq 0},$$T(G)\coloneqq V_{\operatorname*{Lip}(L)}(\mathbb{P}_{n},\mathbb{P}^{G(z)})$ is continuous and $\underset{\mathcal{G}}{\arg\min}V_{\operatorname*{Lip}(L)}(\mathbb{P}^{*},\mathbb{P}^{G(z)})$ is nonempty.
3 From Vanilla to Wasserstein and back
Our subsequent analysis builds on the following equivalence result between the Vanilla GAN distance and the Wasserstein distance with an additional $L^{2}$-penalty term on the discriminator.
Theorem 3.1.
For $L>2$ and $B>0$ we have
$\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,B^{\prime})\\ W(\cdot)>-\log(2-2/L)\end{subarray}}$$\Big{\{}\mathbb{E}[W(X)-W(Y)]-\frac{L(L-1)}{2}\mathbb{E}[W(X)^{2}]\Big{\}}$
$\leq\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})\leq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>-\log(2)\end{subarray}}\Big{\{}\mathbb{E}[W(X)-W(Y)]-\frac{e^{B}}{(2e^{B}-1)^{2}}\mathbb{E}[W(X)^{2}]\Big{\}},$
where $B^{\prime}=\log((1+e^{B})/2).$
Theorem 3.1 reveals that the Vanilla GAN distance is indeed compatible with the Wasserstein distance and will allow us to prove rates of convergence of the Vanilla GAN with respect to the Wasserstein distance. In doing so, we need to consider the consequences of the penalty term. More precisely, we can deduce from Theorem 3.1 that the Vanilla GAN distance is bounded from above and below by the Wasserstein distance or the squared Wasserstein distance, respectively.
Theorem 3.2.
Let $L>2$ and $B\in(0,\infty].$ For two probability measures $\mathbb{P}$ and $\mathbb{Q}$ on $\mathcal{X}$ we have
$\min\Big{(}c_{1}\mathsf{W}_{1}(\mathbb{P},\mathbb{Q}),c_{2}\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})^{2}\Big{)}$ $\leq\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})\leq L\mathsf{W}_{1}(\mathbb{P},\mathbb{Q}),$
where $c_{1}=\frac{1}{2}\frac{\log(2-2/L)}{d^{1/p}}$ and $c_{2}=\frac{1}{2d^{2/p}}\frac{1}{L(L-1)}$, setting $01/p=$ 0 if $p=\infty$.
The assumption that $L>2$ is not very restrictive and takes into account that in practically relevant cases, such as neural network discriminators, the Lipschitz constant is typically quite large. More importantly, we observe a gap between $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})^{2}$ in the lower bound and $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})$ in upper bound when $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})<1$ which is a consequence of the penalty term in Theorem 3.1. The following example indicates that this loss is unavoidable, by restricting the discriminator class to a subset of $\operatorname*{Lip}(L)$.
Example 3.3.
For $\varepsilon,\gamma>0,\gamma+\varepsilon<1$ let $\mathbb{P}=\frac{1}{2}(\delta_{\gamma}+\delta_{\gamma+\varepsilon})$ and $\mathbb{Q}=\frac{1}{2}(\delta_{0}+\delta_{\varepsilon}).$ The Wasserstein distance is then given by
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})=\gamma.$
We consider the Vanilla GAN distance using $L$-Lipschitz affine linear functions as discriminator,$\mathsf{V}_{a\cdot+b}(\mathbb{P},\mathbb{Q}),$ with $a,b\in\mathbb{R}$ and $|a|\leq L.$ Note that the class of affine linear functions can be represented by one layer neural networks (for a definition see Section 5). The optimal $b$ can be calculated explicitly, the optimal $a$ can be determined numerically. Using the optimal slope $a$ and $b$ we obtain for $\gamma<\varepsilon,$ $\varepsilon=\frac{1}{4}$ and $a>16$
$\frac{\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})^{2}}{2}\leq\mathsf{V}_{a\cdot+b}(\mathbb{P},\mathbb{Q})\leq a\cdot\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})^{2}.$
If $\gamma\geq\varepsilon$, then the optimal $a$ is $a=L$ and
$\log(2)\cdot\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})\leq\mathsf{V}_{a\cdot+b}(\mathbb{P},\mathbb{Q})\leq a\cdot\mathsf{W}_{1}(\mathbb{P},\mathbb{Q}).$
See Appendix A for more details on these calculations.
Wasserstein GANs, where the generator is chosen as the empirical risk minimizer of the Wasserstein distance (18), achieve optimal convergence rates (up to logarithmic factors) with respect to the Wasserstein distance as proved by Stéphanovitch et al., (2023). In view of Theorem 3.2 we cannot hope that Vanilla GANs achieve the same rate even if we use a Lipschitz discriminator class. This is in line with the better performance of Wasserstein GANs in practice. However, Theorem 3.2 allows us to study the behavior of Vanilla GANs in settings where the dimension of the latent space is smaller than the dimension of the sample space, a setting that is excluded in all previous work on convergence rates for Vanilla GANs.
4 Oracle inequality for Vanilla GAN in Wasserstein distance
Our aim is to bound the Wasserstein distance between the unknown distribution $\mathbb{P}^{*}$ and the generated distribution $\mathbb{P}^{\hat{G}_{n}(Z)}$ using the empirical risk minimizer $\hat{G}_{n}$ of the Vanilla GAN. The following oracle inequality shows that imposing a Lipschitz constraint on the discriminator class does circumvents the theoretical limitations of Vanilla GANs which is caused by the coupling with the Jensen-Shannon distance. Recall notation (3).
Theorem 4.1.
Let $L>2$ and $B\in(0,\infty].$ For the empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}=\operatorname*{Lip}(L,B)$ we have
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})\leq c\big{[}\inf_{G\in\mathcal{G}}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})\big{]}^{1;1/2}+(1+c)[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})]^{1;1/2},$ (19)
for some constant $c>0$ depending on $d,p$ and $L.$
Note that the discriminator class $\operatorname*{Lip}(L,B)$ admits no finite dimensional parameterization and is therefore not feasible in practice. We will return to this issue in Section 5. The terms in (19) can be interpreted analogously to the interpretation of the bound in Lemma 2.1, but here we have an oracle inequality with respect to the Wasserstein distance. The first term is the approximation error. It is large when $\mathcal{G}$ is not flexible enough to provide a good approximation of $\mathbb{P}^{*}$ by $\mathbb{P}^{G}$ for some $G\in\mathcal{G}$. The second term refers to the stochastic error. With a growing number of observations the empirical measure $\mathbb{P}_{n}$ converges to $\mathbb{P}^{*}$ in Wasserstein distance, see Dudley, (1969), and thus the stochastic error converges to zero. Together with the bounds on $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})$ by Schreuder, (2020) we conclude the following:
Corollary 4.2.
Let $L>2,B\in(0,\infty]$. For some constant $c>0$ the empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}=\operatorname*{Lip}(L,B)$ satisfies
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]\leq$ $\inf_{\begin{subarray}{c}G^{*}\colon\mathcal{Z}\rightarrow\mathcal{X}\end{subarray}}\Big{\{}c[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})]^{1;1/2}+c[\inf_{G\in\mathcal{G}}\|G-G^{*}\|^{1;1/2}_{\infty}]\Big{\}}$
$\qquad\qquad\qquad+c\begin{cases}n^{-1/2d},&d>2,\\ n^{-1/4}(\log n)^{1/2},&d=2,\\ n^{-1/4},&d=1.\end{cases}$
If there is some $G^{*}$ such that *$\mathbb{P}^{*}=\mathbb{P}^{G^{*}(Z)},$ which is commonly assumed in the GAN literature, see e.g. Stéphanovitch et al., (2023),  the first term vanishes and the approximation error is bounded by $\inf_{G\in\mathcal{G}}[\|G-G^{*}\|^{1;1/2}_{\infty}]$. In the bound of the stochastic error we observe the curse of dimensionality: For large dimensions $d$ the rate of convergence $n^{-1/2d}$ deteriorates.
To allow for a dimension reduction setting, we adopt the miss-specified setting from Vardanyan et al., (2023, Theorem 1). In this scenario we can conclude statistical guarantees for Vanilla GANs that are comparable to the results obtained for Wasserstein GANs by Schreuder et al., (2021, Theorem 2). Since Theorem 3.1 does not show equivalence of the Vanilla GAN distance and the Wasserstein distance, we lose a square root in the rate.
Theorem 4.3.
Let $L>2,B\in(0,\infty]$ and $M>0$. The empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}=\operatorname*{Lip}(L,B)$ satisfies
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]$ $\leq\inf_{\begin{subarray}{c}G^{*}\in\operatorname*{Lip}(M,\mathcal{Z})\end{subarray}}\Big{\{}c[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})]^{1;1/2}+c[\inf_{G\in\mathcal{G}}\|G-G^{*}\|^{1;1/2}_{\infty}]\Big{\}}$
$\qquad\qquad\qquad+c\begin{cases}n^{-1/2d^{*}},&d^{*}>2\\ n^{-1/4}(\log n)^{1/2},&d^{*}=2\\ n^{-1/4},&d^{*}=1,\end{cases}$
for some constant $c$ depending on the dimension $d^{*}$ of the latent space $\mathcal{Z}$ and $M$.
The Wasserstein distance $\mathsf{W}_{1}(\mathbb{P}^{G^{*}(Z)},\mathbb{P}^{*})$ now includes also the error due to the dimension reduction while the stochastic error is determined by the potentially much smaller dimension $d^{*}<d$ of the latent space. Compared to Corollary 4.2, the only price for this improvement is the additional Lipschitz restriction on $G^{*}$. We observe a trade-off in the choice of $d^{*}$, since large latent dimensions reduce the approximation error for $\mathbb{P}^{*}$, but increase the stochastic error term.
Theorem 4.3 reveals why Vanilla GANs do perform well in high dimensions. This phenomenon could not be explained in previous work on Vanilla GANs. Belomestny et al., (2021) and Biau et al., (2018) both obtain rates in the Jensen-Shannon distance.
5 Vanilla GANs with network discriminator
In practice, both $\mathcal{G}$ and $\mathcal{D}$ are sets of neural networks. Our conditions on the generator class $\mathcal{G}$ are compactness and good approximation properties of some $G^{*}$ which is chosen such that $\mathbb{P}^{G^{*}}$ mimics $\mathbb{P}^{*}$. Since neural networks have a finite number of weights, and the absolute value of those weights is typically bounded, the compactness assumption is usually satisfied and neural networks enjoy excellent approximation properties, cf. DeVore et al., (2021).
The situation is more challenging for the discriminator class. So far, $\mathcal{D}$ was chosen as the set of Lipschitz functions concatenated to the logistic function. The Lipschitz property is crucial for proof of Theorem 4.1 and thus for all subsequent results.
Controlling the Lipschitz constant while preserving the approximation properties is an area of ongoing research and is far from trivial. Without further restrictions on the class of feedforward networks, the Lipschitz constant would be a term that depends exponentially on the size of the network, see Liang, (2017, Theorem 3.2). Bounding the Lipschitz constant of a neural network is a problem that arises naturally in the implementation of Wasserstein GANs. Arjovsky et al., (2017) use weight clipping to ensure Lipschitz continuity. Later, other approaches such as gradient penalization (see Gulrajani et al., (2017), which was further developed by Wei et al., (2018), Zhou et al., (2019)), Lipschitz penalization (Petzka et al.,, 2017), or spectral penalization (Miyato et al.,, 2018) were introduced and have achieved improved performance in practice.
To extend the theory from the previous section to neural network discriminator classes, we first generalize Theorem 4.1 from $\mathcal{W}=\operatorname*{Lip}(L,B)$ to subsets $\mathcal{W}\subseteq\operatorname*{Lip}(L,B)$. As a result there is an additional approximation error term that accounts for the smaller discriminator class.
Theorem 5.1.
Let $L>2,B\in(0,\infty]$. The empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}\subseteq\operatorname*{Lip}(L,B)$ satisfies for some constant $c>0$ that
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})$ $\leq c\big{[}\inf_{G\in\mathcal{G}}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})\big{]}^{1;1/2}+c\big{[}\inf_{W^{\prime}\in\mathcal{W}}\sup_{W\in\operatorname*{Lip}(L,B)}\|W-W^{\prime}\|_{\infty}\big{]}^{1;1/2}$
$\qquad\qquad\qquad+c\big{[}\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})\big{]}^{1;1/2}.$
To apply this result, we must ensure that the Lipschitz constant of a set of neural networks $\mathcal{W}$ is uniformly bounded by some constant $L$. Adding penalties to the objective function of the optimization problem does not guarantee a fixed bound on the Lipschitz constant. Approaches such as bounds on the spectral or row-sum norm of matrices in feedforward neural networks ensure a bound on the Lipschitz constant, but lead to a loss of expressiveness when considering ReLU networks, even in very simple cases such as the absolute value, see Huster et al., (2019) and Anil et al., (2019). On the other hand, Eckstein, (2020) has shown that one-layer $L$ Lipschitz networks are dense (with respect to the uniform norm) in the set of all $L$ Lipschitz functions on bounded domains. While this implies that the discriminant approximation error converges to zero for growing network architectures, the density statement does not lead to a rate of convergence that depends on the size of the network.
Anil et al., (2019), motivated by Chernodub & Nowicki, (2016), have introduced an adapted activation function, Group Sort, which leads to significantly improved approximation properties of the resulting networks. They show that networks using the Group Sort activation function are dense in the set of Lipschitz functions, but there is no quantitative approximation result. A discussion of the use of Group Sort in the context of Wasserstein GANs can be found in Biau et al., (2021).
To overcome this problem, we would like to approximate not only the optimal discriminating Lipschitz function from the Wasserstein optimization problem in the uniform norm, but also its (weak) derivative. This would allow us to keep the Lipschitz norm of the approximating neural network bounded. For networks with regular activation functions Belomestny et al., (2023) have studied the simultaneous approximation of smooth functions and their derivatives. Gühring et al., (2020) have focused on ReLU networks and have derived quantitative approximation bounds in higher order Hölder and Sobolev spaces. As an intrinsic insight from approximation theory, the regularity of the function being approximated must exceed the regularity order of the norm used to derive approximation bounds. Therefore, we cannot expect to obtain quantitative approximation results for ReLU networks in Lipschitz norm without assuming the continuous differentiability of the approximated function.
Unfortunately, the maximizing function of the Wasserstein optimization problem is in general just Lipschitz continuous. Since we cannot increase the regularity of the target function, we instead relax the Lipschitz assumption of the discriminator in Theorem 5.2 to $\alpha$-Hölder continuity for $\alpha\in(0,1)$. This generalization in the context of Wasserstein GANs has recently been discussed by Stéphanovitch et al., (2023). Recall the definition of the Hölder ball from (5).
Theorem 5.2.
Let $L>2,B>\log(3)$ and $\Gamma>\max(L,2B)$. The empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}\subseteq\mathcal{H}^{\alpha}(\Gamma)$ satisfies for any $G^{*}\in\operatorname*{Lip}(M,\mathcal{Z})$
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]\leq$ $c\big{[}\inf_{G\in\mathcal{G}}\|G^{*}-G\|_{\infty}^{\alpha}\big{]}^{1;1/2}+c\big{[}\mathsf{W}_{1}(\mathbb{P^{*}},\mathbb{P}^{G^{*}(Z)})^{\alpha}\big{]}^{1;1/2}$
$+c\big{[}\inf_{W\in\mathcal{W}}\sup_{W^{*}\in\operatorname*{Lip}(L,B)}\|W-W^{*}\|_{\infty}\big{]}^{1;1/2}$
$+c\begin{cases}n^{-\alpha/2d^{*}},&2\alpha<d^{*}\\ n^{-1/4}(\log n)^{1/2},&2\alpha=d^{*}\\ n^{-1/4},&2\alpha>d^{*}\end{cases}$
for some constant $c$ depending on $d^{*},M$ and $\Gamma$.
It remains to show that there are ReLU networks that satisfy the assumptions of Theorem 5.2. To this end, we build on and extend the approximation results by Gühring et al., (2020).
To fix the notation we give a general definition of feedforward neural networks. Let $d,K,N_{1},...,N_{K}\in\mathbb{N}$. A function $\Phi\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a neural network with $K$ layers and $N_{1}+\dots+N_{K}$ neurons if it results for an argument $x\in\mathbb{R}^{d}$ from the following scheme:
$x_{0}:=x,$ (20)
$x_{k}:=\sigma\left(A_{k}x_{k-1}+b_{k}\right),\quad\text{ for }k=1,\ldots K-1,$
$\Phi(x)=\,$ $x_{K}:=A_{L}x_{K-1}+b_{K},$
where for $k\in\{1,...,K\},$ $A_{k}\in\mathbb{R}^{N_{k}\times N_{k-1}}$ and $b_{k}\in\mathbb{R}^{N_{k}}.$ $\sigma\colon\mathbb{R}\rightarrow\mathbb{R}$ is an element-wise applied arbitrary activation function. The number of nonzero weights of all $A_{k},b_{k}$ is given by $\sum_{j=1}^{K}\left(|A_{j}|_{\mathcal{\ell}^{0}}+|b_{j}|_{\ell^{0}}\right)$. We focus on the ReLU activation function $\sigma(x)=\max(0,x).$
Theorem 5.3.
Let $L,B>0$, and $10<\alpha<$ 1. Then there are constants $C^{\prime},C^{\prime\prime},C^{\prime\prime\prime}>0$ depending on $d,L,\alpha$ and $B$ with the following properties: For any $\varepsilon\in(0,1/2)$ and any $f\in\operatorname*{Lip}(L,B)$, there is a ReLU neural network $\Phi_{\varepsilon}$ with no more than $\lceil C^{\prime}\log_{2}(\varepsilon^{-\frac{1}{1-\alpha}})\rceil$ layers, $\lceil C^{\prime\prime}\varepsilon^{-\frac{d}{1-\alpha}}\log^{2}_{2}(\varepsilon^{-\frac{1}{1-\alpha}})\rceil$ nonzero weights and $\lceil C^{\prime\prime\prime}\varepsilon^{-\frac{d}{1-\alpha}}(\log_{2}^{2}(\varepsilon^{-\frac{1}{1-\alpha}})\vee\log_{2}(\varepsilon^{-\frac{1}{1-\alpha}}))\rceil$ neurons such that
[…]
Since there are many different reasons why controlling the Hölder constant of neural networks is interesting (with stability probably being the most prominent one), Theorem 5.3 is of interest on its own. Combining Theorem 5.2 and Theorem 5.3 with a standard approximation result for the generator approximation error, such as Yarotsky, (2017, Theorem 1), leads to a rate of convergence.
Corollary 5.4.
For $0<\alpha<1,$ $\Gamma>5,$ $M>0,$ $d^{*}>2\alpha$ and $n>2^{\frac{2d^{*}}{\alpha}}$ choose $\mathcal{G}$ as the set of ReLU networks with at most $c\cdot(\log(n)+1)$ layers, $c\cdot n^{-\alpha/2}(\log(n)+1)$ neurons and $c\cdot n^{-\alpha/2}(\log(n)+1)$ nonzero weights and $\mathcal{W}^{\prime}$ as the set of ReLU networks with at most $1+c\cdot\log(n)$ layers and $c\cdot n^{\frac{\alpha}{2(1-\alpha)}}\log^{2}(n)$ nonzero weights and neurons, where $c$ is a constant depending on $d,d^{*},\Gamma$ and $\alpha$. Then the empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}=\mathcal{W}^{\prime}\cap\mathcal{H}^{\alpha}(\Gamma)$ satisfies
$\mathbb{E}\left[\mathrm{W}_{1}\left(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)}\right)\right]\leq c\cdot n^{-\alpha/2d^{*}}+c\big{[}\inf_{G^{*}\in\operatorname*{Lip}(M,\mathcal{Z})}\mathsf{W}_{1}(\mathbb{P^{*}},\mathbb{P}^{G^{*}(Z)})^{\alpha}\big{]}^{1;1/2}.$
This corollary shows that Vanilla GANs with a Hölder regular discriminator class are theoretically advantageous. Since $\alpha$ can be chosen arbitrarily close to one, this reveals why a Lipschitz regularization as implemented for Wasserstein GANs also improves the Vanilla GAN. An empirical confirmation can be found in Zhou et al., (2019).
6 Wasserstein GAN
The same analysis can be applied to Wasserstein-type GANs. The constrained on the Hölder constant can be weakened, as we do not need Theorem 3.2. Note that this does not impact the rate, but the constant. Define the Wasserstein-type GAN with discriminator class $\mathcal{W}$ as
$\mathsf{W}_{\mathcal{W}}(\mathbb{P},\mathbb{Q})=\sup_{W\in\mathcal{W}}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Y)].$
The following theorem shows that by using Hölder continuous ReLU networks as the discriminator class, Wasserstein-type GANs can avoid the curse of dimensionality. Furthermore, this avoids the difficulties arising from the Lipschitz assumption of the neural network, as pointed out by Huang et al., (2022).
Theorem 6.1.
For $0<\alpha<1,$ $\Gamma>2d^{1/p},M>0$ and $d>2\alpha$ and $n>2^{\frac{d}{\alpha}}$ choose $\mathcal{G}$ as the set of ReLU networks with at most $c\cdot(\log(n)+1)$ layers, $c\cdot n^{{\alpha}}(\log(n)+1)$ neurons an nonzero weights and $\mathcal{W}^{\prime}$ as the set of ReLU networks with at most $c\cdot\log(n)$ layers and $c\cdot n^{\frac{\alpha}{(1-\alpha)}}\log^{2}(n)$ nonzero weights and neurons, where $c$ is a constant depending on $d,d^{*},\Gamma$ and $\alpha$ The empirical risk minimizer with $\mathcal{W}=\mathcal{W}^{\prime}\cap\mathcal{H}^{\alpha}(\Gamma)$
$\hat{G}_{n}\in\underset{G\in\mathcal{G}}{\operatorname{argmin}}\;\mathsf{W}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})$
satisfies
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]\leq\min\Big{\{}c\cdot n^{-\frac{\alpha}{d}}+\inf_{G^{*}\mathrm{mb.}}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)}),\;c\cdot n^{-\frac{\alpha}{d^{*}}}+\inf_{G^{*}\in\operatorname{Lip}(M,\mathcal{Z})}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})\Big{\}}.$
Compared to Corollary 5.4 the rate improves to $n^{-\alpha/d^{*}}$ for any $\alpha<1$. This upper bound thus coincides with the lower bound in Tang & Yang, (2022, Theorem 1) up to an arbitrary small polynomial factor. Our rate does not depend exponentially on the number of layers like the results of Liang, (2017), Huang et al., (2022) and we use non-smooth simple ReLU networks compared to smooth ReQU networks in Stéphanovitch et al., (2023) or group sort networks in Biau et al., (2021). In addition, are able to evaluate in the Wasserstein metric rather than the $\mathcal{H}^{\alpha}$ metric.
7 Proofs
7.1 Proof for Section 2
Proof of Lemma 2.1.
Let $X\sim\mathbb{P}^{*},\hat{X}\sim\mathbb{P}_{n}$ and $Z\sim\mathbb{U}$. The symmetry of $\mathcal{W}$ and the Lipschitz continuity of $x\mapsto\log(1+e^{-x})$ yields for any $G\in\mathcal{G}$
$\mathsf{V}_{\mathcal{W}}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}}(Z))$
$=\sup_{W\in\mathcal{W}}\mathbb{E}\Big{[}-\log\Big{(}\frac{1+e^{-W(X)}}{2}\Big{)}+\log\Big{(}\frac{1+e^{-W(\hat{X})}}{2}\Big{)}-\log\Big{(}\frac{1+e^{-W(\hat{X})}}{2}\Big{)}-\log\Big{(}\frac{1+e^{W(\hat{G}_{n}(Z))}}{2}\Big{)}\Big{]}$
$\leq\sup_{W\in\mathcal{W}}\mathbb{E}\big{[}-\log(1+e^{-W(X)})+\log(1+e^{-W(\hat{X})})\big{]}+\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$
$\leq\sup_{W\in\mathcal{W}}\mathbb{E}\big{[}-\log(1+e^{-W(X)})+\log(1+e^{-W(\hat{X})})\big{]}+\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})$
$=\sup_{W\in\mathcal{W}}\mathbb{E}[-\log(1+e^{-W(X)})+\log(1+e^{-W(\hat{X})})]$
$\quad+\sup_{W\in\mathcal{W}}\mathbb{E}[-\log(1+e^{-W(\hat{X})})+\log(1+e^{-W(X)})]+\mathsf{V}_{\mathcal{W}}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})$
$\leq 2\sup_{W\in\operatorname*{Lip}(1)\circ\mathcal{W}}\mathbb{E}[W(X)-W(\hat{X})]+\mathsf{V}_{\mathcal{W}}(\mathbb{P}^{*},\mathbb{P}^{G(Z)}).$
The bound for $\hat{G}_{n}$ from (10) follows since $G\in\mathcal{G}$ was arbitrary. ∎
Proof of Lemma 2.2.
Let $(G_{n})_{n\in\mathbb{N}}\in\mathcal{G}$ be a sequence that converges to $G\in\mathcal{G}.$ If $\mathsf{V}_{\operatorname*{Lip}(L)}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})\geq\mathsf{V}_{\operatorname*{Lip}(L)}(\mathbb{P}^{*},\mathbb{P}^{G_{n}(Z)})$, then
$\mathsf{V}_{\operatorname*{Lip}(L)}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})-\mathsf{V}_{\operatorname*{Lip}(L)}(\mathbb{P}^{*},\mathbb{P}^{G_{n}(Z)})$
$\leq\sup_{W\in\operatorname*{Lip}(L)}\mathbb{E}\Big{[}\log\Big{(}\frac{1+e^{-W(G_{n}(Z))}}{2}\Big{)}-\log\Big{(}\frac{1+e^{-W(G(Z))}}{2}\Big{)}\Big{]}$
$\leq\sup_{W\in\operatorname*{Lip}(L)}\mathbb{E}\Big{[}W(G_{n}(Z))-W(G(Z))\Big{]}$
$\leq L\|G_{n}-G\|_{\infty}.$
The case $\mathsf{V}_{\operatorname*{Lip}(L)}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})<\mathsf{V}_{\operatorname*{Lip}(L)}(\mathbb{P}^{*},\mathbb{P}^{G_{n}(Z)})$ can be bounded analogously. Therefore, $T$ is continuous and there is at least one minimizer if $\mathcal{G}$ is compact. ∎
7.2 Proofs for Section 3
Before we prove the main results from Section 3 we require an auxiliary lemma:
Lemma 7.1.
For $X\sim\mathbb{P}$ and $Y\sim\mathbb{Q}$ and an arbitrary set of measurable functions $\mathcal{W}$ we have that
$\mathsf{V}_{\mathcal{W}}(\mathbb{P},\mathbb{Q})\leq\sup_{W\in\mathcal{W}}\mathbb{E}[-\log\big{(}1+e^{-W(X)}\big{)}+\log\big{(}1+e^{-W(Y)}\big{)}].$
Proof.
Since
$\log(1+e^{x})+\log(1+e^{-x})\geq\log(4)\quad\text{ for all }x\in\mathbb{R},$
we can bound
$\sup_{W\in\mathcal{W}}$ $\mathbb{E}\Big{[}-\log\big{(}1+e^{-W(X)}\big{)}-\log\big{(}1+e^{W(Y)}\big{)}\Big{]}$
$=\sup_{W\in\mathcal{W}}\mathbb{E}[-\log\big{(}1+e^{-W(X)}\big{)}+\log\big{(}1+e^{-W(Y)}\big{)}-\log\big{(}1+e^{-W(Y)}\big{)}-\log\big{(}1+e^{W(Y)}\big{)}]$
$\leq\sup_{W\in\mathcal{W}}\mathbb{E}[-\log\big{(}1+e^{-W(X)}\big{)}+\log\big{(}1+e^{-W(Y)}\big{)}]\quad-\inf_{W\in\mathcal{W}}\mathbb{E}[\log\big{(}1+e^{-W(Y)}\big{)}+\log\big{(}1+e^{W(Y)}\big{)}]$
$\leq\sup_{W\in\mathcal{W}}\mathbb{E}[-\log\big{(}1+e^{-W(X)}\big{)}+\log\big{(}1+e^{-W(Y)}\big{)}]-\log(4).\qed$
Proof of Theorem 3.1.
Defining
$\psi\colon\mathbb{R}\rightarrow\mathbb{R},\ \psi(x)\coloneqq-\log\Big{(}\frac{1+e^{-x}}{2}\Big{)},$
we can rewrite
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})=\sup_{W\in\operatorname*{Lip}(L,B)}\mathbb{E}[\psi(W(X))+\psi(-W(Y))].$
The function $f\colon[-\log(2-2/L),\infty)\rightarrow\mathbb{R},$ $\ f(x)=\log(2e^{x}-1)$ is bijective and Lipschitz continuous with Lipschitz constant $L$ and satisfies $\psi(-f(x))=x$ for all $x\geq-\log(2-2/L)$. Therefore, we obtain a lower bound
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})$ $\geq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,\log((1+e^{B})/2))\\ W(\cdot)\geq-\log(2-2/L)\end{subarray}}\mathbb{E}[\psi(f(W(X)))+\psi(-f(W(Y)))]$
$=\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,B^{\prime})\\ W(\cdot)\geq-\log(2-2/L)\end{subarray}}\mathbb{E}[\psi(f(W(X)))-W(Y)].$
Since $f^{-1}\in\operatorname*{Lip}(1,\mathbb{R})$, we can estimate $\mathsf{V}_{\operatorname*{Lip}(L,B)}$ from above by
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})$ $=\sup_{W\in\operatorname*{Lip}(L,B)}\mathbb{E}[\psi(f(f^{-1}(W(X))))+\psi(-f(f^{-1}(W(Y))))]$
$\leq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>\log(2)\end{subarray}}\mathbb{E}[\psi(f(W(X)))+\psi(-f(W(Y)))]$
$=\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>\log(2)\end{subarray}}\mathbb{E}[\psi(f(W(X)))-W(Y)].$
A Taylor approximation at zero of the function $\psi\circ f(x)=\log(2-e^{-x})$ yields that for every $x\in(-\log(2),\infty)$ there exists a $\xi$ between $x$ and 00 such that
$\psi\circ f(x)=x-\frac{e^{\xi}}{(2e^{\xi}-1)^{2}}x^{2}.$
For the lower bound, we thus conclude
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})\geq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,B^{\prime})\\ W(\cdot)\geq-\log(2-2/L)\end{subarray}}\mathbb{E}[W(X)-W(Y)]-\frac{L(L-1)}{2}\mathbb{E}[W(X)^{2}].$
For the upper bound, we get
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})\leq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>\log(2)\end{subarray}}\mathbb{E}[W(X)-W(Y)]-\frac{e^{B}}{(2e^{B}-1)^{2}}\mathbb{E}[W(X)^{2}].\qed$
Note that, using the function $g\colon(-\infty,\log(2-2/L))\rightarrow\mathbb{R},\;g(x)=-\log(2e^{-x}-1),$ we obtain lower and upper bounds with a penalty term depending on $\mathbb{E}[W(Y)^{2}]$ instead of $\mathbb{E}[W(X)^{2}]$.
Proof of Theorem 3.2.
We prove the lower bound first. Theorem 3.1 yields
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})$ $\geq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,B^{\prime})\\ W(\cdot)>-\log(2-2/L)\end{subarray}}\mathbb{E}[W(X)-W(Y)]-\frac{L(L-1)}{2}\mathbb{E}[W(X)^{2}]$
$\geq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,\log(2-2/L))\end{subarray}}\mathbb{E}[W(X)-W(Y)]-\frac{L(L-1)}{2}\mathbb{E}[W(X)^{2}].$
Let $W^{*}\in\underset{\begin{subarray}{c}W\in\operatorname*{Lip}(1,\log(2-2/L))\end{subarray}}{\arg\max}\mathbb{E}[W(X)-W(Y)].$ Then $\delta W^{*}\in\operatorname*{Lip}(1,\log(2-2/L))$ for all $\delta\in(0,1]$ and we can conclude
$\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,\log(2-2/L))\end{subarray}}\mathbb{E}[W(X)$ $-W(Y)]-\frac{L(L-1)}{2}\mathbb{E}[W(X)^{2}]$
$\geq\sup_{\delta\in(0,1]}\Big{\{}\mathbb{E}[\delta W^{*}(X)-\delta W^{*}(Y)]-\frac{L(L-1)}{2}\mathbb{E}[(\delta W^{*}(X))^{2}]\Big{\}}$
$=\sup_{\delta\in(0,1]}\Big{\{}\delta\mathbb{E}[W^{*}(X)-W^{*}(Y)]-\delta^{2}\frac{L(L-1)}{2}\mathbb{E}[(W^{*}(X))^{2}]\Big{\}},$
which is independent from $B$. In case $\Delta:=\mathbb{E}[W^{*}(X)-W^{*}(Y)]<L(L-1)\mathbb{E}[W^{*}(X)^{2}]$ we have for $\delta=\frac{\Delta}{\mathbb{E}[W^{*}(X)^{2}]L(L-1)}\in(0,1)$
$\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,\log(2-2/L))\end{subarray}}\mathbb{E}[W(X)-W(Y)]-\frac{L(L-1)}{2}\mathbb{E}[W(X)^{2}]$
$\qquad\qquad\qquad\geq\frac{\Delta^{2}}{\mathbb{E}[W^{*}(X)^{2}]L(L-1)}-\frac{\Delta^{2}}{2\mathbb{E}[W^{*}(X)^{2}]L(L-1)}$
$\qquad\qquad\qquad=\frac{\Delta^{2}}{2\mathbb{E}[W^{*}(X)^{2}]L(L-1)}$
$\qquad\qquad\qquad\geq\frac{\Delta^{2}}{2\log(2-2/L)^{2}L(L-1)},$
where we used $|W^{*}(x)|\leq\log(2-2/L)$ in the last step. In case $\Delta\geq L(L-1)\mathbb{E}[W^{*}(X)^{2}]$ we obtain
$\mathbb{E}[W^{*}(X)-W^{*}(Y)]-\frac{L(L-1)}{2}\mathbb{E}[W^{*}(X)^{2}]$ $\geq\frac{1}{2}\mathbb{E}[W^{*}(X)-W^{*}(Y)].$
Using
$\Delta$ $=\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,\log(2-2/L))\end{subarray}}\mathbb{E}[W(X)-W(Y)]$
$\geq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(\log(2-2/L))d^{-1/p},\infty)\end{subarray}}\mathbb{E}[W(X)-W(Y)]=\frac{\log(2-2/L))}{d^{1/p}}\mathsf{W}_{1}(\mathbb{P},\mathbb{Q}),$
we conclude the claimed lower bound for
$c_{1}=\frac{1}{2}\frac{\log(2-2/L)}{d^{\frac{1}{p}}},\quad c_{2}=\frac{1}{2d^{\frac{2}{p}}L(L-1)}.$
For the upper bound we use Lemma 7.1 with $\mathcal{W}=\operatorname*{Lip}(L).$ Since for $W\in\operatorname*{Lip}(L)$ the function $-\log\big{(}1+e^{-W(\cdot)}\big{)}\in\operatorname*{Lip}(L)$ we conclude
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})$ $\leq\sup_{W\in\operatorname*{Lip}(L)}\mathbb{E}[\psi(W(X))+\psi(W(Y))]$
$\leq\sup_{W\in\operatorname*{Lip}(L)}\mathbb{E}[-\log\big{(}1+e^{-W(X)}\big{)}+\log\big{(}1+e^{-W(Y)}\big{)}]$
$\leq\sup_{W\in\operatorname*{Lip}(L)}\mathbb{E}[W(X)-W(Y)]$
$=L\sup_{W\in\operatorname*{Lip}(1)}\mathbb{E}[W(X)-W(Y)].\qed$
7.3 Proofs for Section 4
Proof of Theorem 4.1.
Using Theorem 3.2 and the triangle inequality for the Wasserstein distance, we deduce for every $G\in\mathcal{G}$ and $c=\max(c_{1}^{-1},c_{2}^{-1/2})$ that
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})$ $\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+c\big{[}\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})\big{]}^{1;1/2}$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+c\big{[}\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})\big{]}^{1;1/2}$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+cL\big{[}\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})\big{]}^{1;1/2}$
$\leq(1+cL)\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})\big{]}^{1;1/2}+cL\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})\big{]}^{1;1/2}.$
As $G\in\mathcal{G}$ was arbitrary, we can choose the infimum over $\mathcal{G}.$
Proof of Corollary 4.2.
For every measurable $G^{*}\colon\mathcal{Z}\rightarrow\mathcal{X}$ and any $G\in\mathcal{G}$ we have
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})$ $\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})+\mathsf{W}_{1}(\mathbb{P}^{G^{*}(Z)},\mathbb{P}^{G(Z)})$
$=\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})+\sup_{W\in\operatorname*{Lip}(1)}\mathbb{E}[W(G^{*}(Z))-W(G(Z))]$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})+\mathbb{E}[|G^{*}(Z)-G(Z)|_{p}]$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})+\|G^{*}-G\|_{\infty}.$
Since $G^{*}$ was arbitrary, Theorem 4.1 yields for some constant $c$
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})]\leq$ $c\cdot\mathbb{E}[\max(\sqrt{\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})},\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*}))]$
$+c\cdot\inf_{\begin{subarray}{c}G^{*}\colon\mathcal{Z}\rightarrow\mathcal{X}\end{subarray}}\Big{\{}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})]^{1;1/2}+[\inf_{G\in\mathcal{G}}\|G-G^{*}\|^{1;1/2}_{\infty}]\Big{\}}$
Using Jensen’s inequality, we can bound the stochastic error term by
$\mathbb{E}[\max(\sqrt{\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})},\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*}))]$ $\leq\mathbb{E}[\sqrt{\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})}]+\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})]$
$\leq\sqrt{\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})]}+\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})].$
From Schreuder, (2020, Theorem 4) we know
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})]\leq c^{\prime}\begin{cases}n^{-1/d},&d>2\\ n^{-1/2}\log(n),&d=2\\ n^{-1/2},&d=1.\end{cases}$
where $c$ depends only on $d$. Since $(\log n)/\sqrt{n}\leq 1$, we conclude
$\sqrt{\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})]}+\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})]\leq 2c^{\prime}\begin{cases}n^{-1/2d},&d>2\\ n^{-1/4}(\log n)^{1/2},&d=2\\ n^{-1/4},&d=1.\end{cases}$
Proof of Theorem 4.3.
With the same reasoning as in the proof of Corollary 4.2, there exists some $c$ such that for any measurable $G^{*}\colon\mathcal{Z}\to\mathcal{X}$ and any $G\in\mathcal{G}$
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G(Z)})]$ $\leq c\big{(}\sqrt{\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})]}+\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})]+[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})]^{1;1/2}+[\inf_{G\in\mathcal{G}}\|G^{*}-G\|_{\infty}]^{1,1/2}\big{)}$
By the triangle inequality
$\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})\leq\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{G^{*}(Z)})+\mathsf{W}_{1}(\mathbb{P}^{G^{*}(Z)},\mathbb{P}^{*}).$
Let $Z_{i}\sim\mathbb{U}$ be i.i.d. random variables and denote the corresponding empirical measure by $\mathbb{U}_{n}$. For $G^{*}\in\operatorname*{Lip}(M,\mathcal{Z})$ we can then bound the first term by
$\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{G^{*}(Z)})$ $=\sup_{W\in\operatorname*{Lip}(1)}\frac{1}{n}\sum_{i=1}^{n}W(X_{i})-\mathbb{E}[W\circ G^{*}(Z)]$
$\leq\sup_{W\in\operatorname*{Lip}(1)}\frac{1}{n}\sum_{i=1}^{n}|W(X_{i})-W\circ G^{*}(Z_{i})|+\sup_{W\in\operatorname*{Lip}(1)}\frac{1}{n}\sum_{i=1}^{n}W\circ G^{*}(Z_{i})-\mathbb{E}[W\circ G^{*}(Z)]$
$\leq\frac{1}{n}\sum_{i=1}^{n}|X_{i}-G^{*}(Z_{i})|_{p}+\sup_{f\in\operatorname*{Lip}(M)}\frac{1}{n}\sum_{i=1}^{n}f(Z_{i})-\mathbb{E}[f(Z)]$ (21)
$=\frac{1}{n}\sum_{i=1}^{n}|X_{i}-G^{*}(Z_{i})|_{p}+M\cdot\mathsf{W}_{1}(\mathbb{U}_{n},\mathbb{U})$
Hence,
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{G^{*}(Z)})\big{]}\leq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[|X_{i}-G^{*}(Z_{i})|_{p}]+M\cdot\mathbb{E}[W_{1}(\mathbb{U}_{n},\mathbb{U})].$
Note that $\mathbb{E}[|X_{i}-G^{*}(Z_{i})|_{p}]=\mathsf{W}_{1}(\mathbb{P}^{G^{*}(Z)},\mathbb{P}^{*})$ by the duality formula of $\mathsf{W}_{1}$ used in this work, see Villani, (2008, Definition 6.2 and Remark 6.5). For $\mathbb{E}[W_{1}(\mathbb{U}_{n},\mathbb{U})]$, we can exploit the convergence rate for the empirical distribution as in Corollary 4.2, but now in the $d^{*}$ dimensional latent space $\mathcal{Z}$. Therefore, there exists a $c^{\prime}$ such that
$\sqrt{\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})]}+\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})]\leq c^{\prime}[\mathsf{W}_{1}(\mathbb{P}^{G^{*}(Z)},\mathbb{P}^{*})]^{1;1/2}+c^{\prime}\begin{cases}n^{-1/2d^{*}},&d^{*}>2\\ n^{-1/4}(\log n)^{1/2},&d^{*}=2\\ n^{-1/4},&d^{*}=1.\end{cases}$
7.4 Proofs of Theorem 5.1 and Theorem 5.2
Proof of Theorem 5.1.
First, we verify that for any two nonempty sets $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ we have
$\mathsf{V}_{\mathcal{W}_{1}}(\mathbb{P},\mathbb{Q})\leq\mathsf{V}_{\mathcal{W}_{2}}(\mathbb{P},\mathbb{Q})+2\inf_{W\in\mathcal{W}_{2}}\sup_{W^{*}\in\mathcal{W}_{1}}\|W-W^{*}\|_{\infty}.$ (22)
Indeed, the difference $\mathsf{V}_{\mathcal{W}_{1}}(\mathbb{P},\mathbb{Q})-\mathsf{V}_{\mathcal{W}_{2}}(\mathbb{P},\mathbb{Q})$ is bounded by
$\inf_{W\in\mathcal{W}_{2}}\sup_{W^{*}\in\mathcal{W}_{1}}\Big{\{}\mathbb{E}\Big{[}-\log\Big{(}\frac{1+e^{-W^{*}(X)}}{2}\Big{)}-\log\Big{(}\frac{1+e^{W^{*}(Y)}}{2}\Big{)}\Big{]}$
$\qquad\qquad\qquad\quad-\mathbb{E}\Big{[}-\log\Big{(}\frac{1+e^{-W(X)}}{2}\Big{)}-\log\Big{(}\frac{1+e^{W(Y)}}{2}\Big{)}\Big{]}\Big{\}}$
$\leq$ $\inf_{W\in\mathcal{W}_{2}}\sup_{W^{*}\in\mathcal{W}_{1}}\Big{\{}\mathbb{E}\Big{[}\Big{|}-\log\Big{(}\frac{1+e^{-W^{*}(X)}}{2}\Big{)}+\log\Big{(}\frac{1+e^{-W(X)}}{2}\Big{)}\Big{|}\Big{]}$
$\qquad\qquad\qquad\quad+\mathbb{E}\Big{[}\Big{|}-\log\Big{(}\frac{1+e^{W^{*}(Y)}}{2}\Big{)}+\log\Big{(}\frac{1+e^{W(Y)}}{2}\Big{)}\Big{|}\Big{]}\Big{\}}$
$\leq\inf_{W\in\mathcal{W}_{2}}\sup_{W^{*}\in\mathcal{W}_{1}}\big{\{}\mathbb{E}[|W^{*}(X)-W(X)|]+\mathbb{E}[|W^{*}(Y)-W(Y)|]\big{\}}$
$\leq 2\inf_{W\in\mathcal{W}_{2}}\sup_{W^{*}\in\mathcal{W}_{1}}\|W^{*}-W\|_{\infty},$
due to Lipschitz continuity of $x\mapsto-\log((1+e^{x})/2)$. From (22) we deduce for $\mathcal{W}\subset\operatorname*{Lip}(L,B)$
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})\leq\mathsf{V}_{\mathcal{W}}(\mathbb{P},\mathbb{Q})+2\inf_{W^{\prime}\in\mathcal{W}}\sup_{W\in\operatorname*{Lip}(L,B)}\|W-W^{\prime}\|_{\infty}.$
We abbreviate $\Delta_{\mathcal{W}}\coloneqq\inf_{W^{\prime}\in\mathcal{W}}\sup_{W\in\operatorname*{Lip}(L,B)}\|W-W^{\prime}\|_{\infty}.$ Now we can proceed as in Theorem 4.1. In particular, it is sufficient to bound $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)}).$ Due to Theorem 3.2 there is some constant $c>0$ such that for every $G\in\mathcal{G}$
$\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$ $\leq c[\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})]^{1;1/2}$
$\leq c[\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})+2\Delta_{\mathcal{W}}]^{1;1/2}$
$\leq c[\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})]^{1;1/2}+2c[\Delta_{\mathcal{W}}]^{1;1/2}$
$\leq c[\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})]^{1;1/2}+2c[\Delta_{\mathcal{W}}]^{1;1/2}$
Because $\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})\leq\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})$ due to $\mathcal{W}\subset\operatorname*{Lip}(L,B)$, the rest of the proof is identical to the proof of Theorem 4.1. ∎
Proof of Theorem 5.2.
Since $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})$ can be estimated as in Theorem 4.3, we only need to bound $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$. For $L=\Gamma,B=\Gamma/2$ we have $\operatorname*{Lip}(L,B)\subset\mathcal{H}^{\alpha}(\Gamma)$, $\alpha\in(0,1)$, and the assumptions of Theorem 3.2 are satisfied. Therefore for every $\alpha\in(0,1)$
$\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$ $\leq c\big{[}\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})\big{]}^{1;1/2}\leq c\big{[}\mathsf{V}_{\mathcal{H}^{\alpha}(\Gamma)}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})\big{]}^{1;1/2}.$
Now, (22) yields
$\mathsf{V}_{\mathcal{H}^{\alpha}(\Gamma)}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})\leq\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})+2\Delta_{\mathcal{W}}\qquad\text{for}\qquad\Delta_{\mathcal{W}}:=\inf_{W\in\mathcal{W}}\sup_{W^{*}\in\mathcal{H}^{\alpha}(\Gamma)}\|W^{*}-W\|_{\infty}.$
Using that $\hat{G}_{n}$ is the empirical risk minimizer and $\mathcal{W}\subseteq\mathcal{H}^{\alpha}(\Gamma)$, we thus have
$\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$ $\leq c\big{[}\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})\big{]}^{1;1/2}+c[\Delta_{\mathcal{W}}]^{1;1/2}$
$\leq c\big{[}\mathsf{V}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})\big{]}^{1;1/2}+c[\Delta_{\mathcal{W}}]^{1;1/2}$
$\leq c\big{[}\mathsf{V}_{\mathcal{H}^{\alpha}(\Gamma)}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})\big{]}^{1;1/2}+c[\Delta_{\mathcal{W}}]^{1;1/2}$
To bound the first term, we apply Lemma 7.1 and $\{-\log(1+e^{-W(\cdot)})\mid W\in\mathcal{H}^{\alpha}(\Gamma)\}\subset\mathcal{H}^{\alpha}(\Gamma)$ to obtain
$\mathsf{V}_{\mathcal{H}^{\alpha}(\Gamma)}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})$ $\leq\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}\hat{X}\sim\mathbb{P}_{n}\end{subarray}}[-\log\Big{(}1+e^{-W(\hat{X})}\Big{)}+\log\Big{(}1+e^{-W(G(Z))}\Big{)}]$
$\leq\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}\hat{X}\sim\mathbb{P}_{n}\\ \end{subarray}}[W(\hat{X})-W(G(Z))]$
$\leq\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}\hat{X}\sim\mathbb{P}_{n}\end{subarray}}[W(\hat{X})-W(X)]+\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}[W(X)-W(G(Z))].$
For the second term we have by Hölder continuity, Jensens inequality and the duality formula of $\mathsf{W}_{1}$ as used in the proof of Theorem 4.3 that
$\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}[W(X)-W(G(Z))]$ $\leq\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}[|W(X)-W(G^{*}(Z))|]+\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}[|W(G^{*}(Z))-W(G(Z))|]$
$\leq\Gamma\mathbb{E}[|X-G^{*}(Z)|_{p}^{\alpha}]+\Gamma\|G^{*}-G\|_{\infty}^{\alpha}$
$\leq\Gamma\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})^{\alpha}+\Gamma\|G^{*}-G\|_{\infty}^{\alpha}.$
Hence, we have for any $G\in\mathcal{G}$ and any measurable $G^{*}\colon\mathcal{Z}\to\mathcal{X}$ for some constant $c>0$
$\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})\leq$ $c\big{[}\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}\hat{X}\sim\mathbb{P}_{n}\end{subarray}}[W(\hat{X})-W(X)]\big{]}^{1;1/2}$
$\qquad+c\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})^{\alpha}+\|G^{*}-G\|_{\infty}^{\alpha}\big{]}^{1;1/2}+c[\Delta_{\mathcal{W}}]^{1;1/2}.$
For the remaining stochastic error term, we first note that
$\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}\hat{X}\sim\mathbb{P}_{n}\end{subarray}}[W(\hat{X})-W(X)]$ $\leq\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}X_{n}\sim\mathbb{P}_{n}\end{subarray}}[W(X_{n})-W(G^{*}(Z))]$
$\qquad+\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}[W(G^{*}(Z))-W(X)]$
$\leq\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}X_{n}\sim\mathbb{P}_{n}\end{subarray}}[W(X_{n})-W(G^{*}(Z))]+\Gamma\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})^{\alpha}$
and as in (21) together with Schreuder, (2020, Theorem 4) we obtain
$\mathbb{E}\Big{[}\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}X_{n}\sim\mathbb{P}_{n}\end{subarray}}[W(X_{n})-W(G^{*}(Z))]\Big{]}$ $\leq\mathbb{E}\Big{[}\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}|X-G^{*}(Z)|_{p}^{\alpha}\Big{]}$
$\qquad+\mathbb{E}\Big{[}\sup_{f\in\mathcal{H}^{\alpha}(M\cdot\Gamma)}\frac{1}{n}\sum_{i=1}^{n}f(Z_{i})-\mathbb{E}[f(Z)]\Big{]}$
$\leq c\mathsf{W}_{1}(\mathbb{P^{*}},\mathbb{P}^{G^{*}(Z)})^{\alpha}+c\begin{cases}n^{-\alpha/d^{*}},&2\alpha<d^{*},\\ n^{-1/2}\ln(n),&2\alpha=d^{*},\\ n^{-1/2},&2\alpha>d^{*}.\end{cases}$
For the expectation of the first term we use Jensen’s inequality
$\mathbb{E}[|X_{i}-G^{*}(Z_{i})|_{p}^{\alpha}]\leq\mathbb{E}[|X_{i}-G^{*}(Z_{i})|_{p}]^{\alpha}=\mathsf{W}_{1}(\mathbb{P^{*}},\mathbb{P}^{G^{*}(Z)})^{\alpha}.\qed$
7.5 Proof of Theorem 5.3
To prove Theorem 5.3 some additional notation is required. The set of locally integrable functions is given by
$L_{\text{loc}}^{1}(\Omega)\coloneqq\{f\colon\Omega\rightarrow\mathbb{R}\Big{|}\int_{K}|f(x)|\;\mathrm{d}x<\infty,\;\text{for all compact }K\subset\Omega^{\circ}\}.$
A function $f\in L_{\text{loc}}^{1}(\Omega)$ has a weak derivative, $D_{w}^{\alpha}f$, provided there exists a function $g\in L_{\text{loc}}^{1}(\Omega)$ such that
[…]
If such a $g$ exists, we define $D_{w}^{\alpha}f\coloneqq g.$ For $f\in L_{\text{loc}}^{1}(\Omega)$ and $k\in\mathbb{N}_{0}$ the Sobolev norm is
$\|f\|_{W^{k,\infty}(\Omega)}:=\max_{|\alpha|\leq k}\left\|D_{w}^{\alpha}f\right\|_{\infty,\Omega}.$
The Sobolev space $W^{k,\infty}(\Omega):=\{f\in L_{\text{loc}}^{1}(\Omega):\|f\|_{W_{p}^{k}(\Omega)}<\infty\}$ is a Banach space (Brenner & Scott,, 2008, Theorem 1.3.2). For $f\in W^{k,\infty}(\Omega)$, define the Sobolev semi norm by
$|f|_{W^{k,\infty}(\Omega)}\coloneqq\max_{|\alpha|=k}\left\|D_{w}^{\alpha}f\right\|_{{\infty},\Omega}.$
Note that $\operatorname*{Lip}(L,B,\Omega)\subset W^{1,\infty}(\Omega)$, since $\|f\|_{W^{1,\infty}}\leq\max(L,B)$ for any $f\in\operatorname*{Lip}(L,B,\Omega)$. For two normed spaces $(A,\|\cdot\|_{A}),(B,\|\cdot\|_{B})$ we denote the operator norm of a linear operator $T\colon A\rightarrow B$ by
$\|T\|:=\sup\big{\{}\|Tx\|_{B}\;|\;x\in A,\|x\|_{A}\leq 1\big{\}}.$
Theorem 5.3 is very close to Theorem 4.1 by Gühring et al., (2020), which however applies only to functions $f$ which are at least twice (weakly) differentiable. Our proof can thus build on numerous auxiliary results and arguments from Gühring et al., (2020). We basically keep the proof structure of Gühring et al., (2020) which in turn relies on Yarotsky, (2017).
Let $d,N\in\mathbb{N}.$ For $m\in\{0,...,N\}^{d},$ define the functions $\phi_{m}\colon\mathbb{R}^{d}\rightarrow\mathbb{R},$
$\phi_{m}(x)=\prod_{\ell=1}^{d}\psi\Big{(}3N\Big{(}x_{\ell}-\frac{m_{\ell}}{N}\Big{)}\Big{)},\quad\text{where}\quad\psi(x)=\begin{cases}1,&|x|<1,\\ 0,&|x|>2,\\ 2-|x|,&1\leq|x|\leq 2\end{cases}$
By definition, we have $\|\phi_{m}\|_{\infty}=1$ for all $m$ and
$\operatorname{supp}\phi_{m}\subset\left\{x:\left|x_{k}-\frac{m_{k}}{N}\right|<\frac{1}{N}\forall k\right\}\eqqcolon B_{\frac{1}{N},|\cdot|_{\infty}}(\frac{m}{N}).$ (23)
Gühring et al., (2020, Lemma C.3 (iv)) have verified that $\|\phi_{m}\|_{W^{1,\infty}(\mathbb{R}^{d})}\leq cN$ for some constant $c>0$.
A direct consequence of Lemma 2.11, Lemma C.3, Lemma C.5 and Lemma C.6 by Gühring et al., (2020) is the following approximation result for the localizing functions $\phi_{m}$ via ReLU networks:
Lemma 7.2.
For any $\varepsilon\in(0,1/2)$ and any $m\in\{0,...,N\}^{d}$ there is a network $\Psi_{\varepsilon}$ with ReLU activation function, not more than $C_{1}\log_{2}(\varepsilon^{-1})$ layers and no more than $C_{2}(N+1)^{d}\log_{2}^{2}(\varepsilon^{-1})$ nonzero weights and no more than neurons $C_{3}(N+1)^{d}(\log_{2}^{2}(\varepsilon^{-1})\vee\log_{2}(\varepsilon^{-1}))$ such that for $k\in\{0,1\}$
$\|\Psi_{\varepsilon}-\phi_{m}\|_{W^{k,\infty}}\leq cN^{k}\varepsilon,$
where $C_{1},C_{2},C_{3}$ and $c$ are constants independent of $m$ and $\varepsilon$. Additionally,
$\phi_{m}(x)=0\Longrightarrow\Psi_{\varepsilon}(x)=0,$
and therefore $\operatorname{supp}\Psi_{\varepsilon}\subset B_{\frac{1}{N},|\cdot|_{\infty}}(\frac{m}{N}).$
Next we approximate a bounded Lipschitz function using linear combinations of the set $\{\phi_{m}\colon m\in\{1,...,N\}^{d}\}$. The approximation error will be measured in the Hölder norm from (4).
Lemma 7.3.
Let $10<\alpha<$ 1 . There exists a constant $C_{1}>0$ such that for any $f\in W^{1,\infty}((0,1)^{d})$ there are constants $c_{f,m}$ for $m\in\{0,...,N\}^{d}$ such that
$\Big{\|}f-\sum_{m\in\{0,...,N\}^{d}}c_{f,m}\phi_{m}\Big{\|}_{\mathcal{H}^{\alpha}}\leq C_{1}\Big{(}\frac{1}{N}\Big{)}^{1-\alpha}\|f\|_{W^{1,\infty}}.$
The coefficients satisfy for a $C_{2}>0$
$|c_{f,m}|\leq C_{2}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})},$
where $\Omega_{m,N}\coloneqq B_{\frac{1}{N},|\cdot|_{\infty}}(\frac{m}{N})$ and $\tilde{f}\in W^{1,\infty}(\mathbb{R})$ is an extension of $f.$
Proof.
Let $E\colon W^{1,\infty}((0,1)^{d})\rightarrow W^{1,\infty}(\mathbb{R})$ be the continuous linear extension operator from (Stein,, 1970, Theorem 5) and set $\tilde{f}\coloneqq Ef.$ As $E$ is continuous there exists a $C_{E}>0$ such that
$\|\tilde{f}\|_{W^{1,\infty}(\mathbb{R}^{d})}\leq C_{E}\|f\|_{W^{1,\infty}}.$
Step 1 (Choice of $c_{f,m}$): For each $m\in\{0,\ldots,N\}^{d}$ we define
$c_{f,m}=\int_{B_{m,N}}\tilde{f}(y)\rho(y)\ \mathrm{d}y\text{ for }B_{m,N}:=B_{\frac{3}{4N},\mid\cdot\mid}\Big{(}\frac{m}{N}\Big{)}$
and an arbitrary cut-off function $\rho$ supported in $B_{m,N}$, i.e.
[…]
Then
$|c_{m,f}|=\Big{|}\int_{B_{m,N}}\tilde{f}(y)\rho(y)\ \mathrm{d}y\Big{|}\leq\|\tilde{f}\|_{\infty,\Omega_{m,N}}\int_{B_{m,N}}\rho(y)\ \mathrm{d}y=\|\tilde{f}\|_{{\infty},\Omega_{m,N}}\leq C_{E}\|f\|_{W^{1,\infty}(\Omega_{m,N})}.$
Step 2 (Local estimates in $\|\cdot\|_{W^{k,p}}$): The coefficients $c_{m,f}$ are the averaged Taylor polynomials in the sense of Brenner & Scott, (2008, Definition 4.1.3) of order $1$ averaged over $B_{m,N}.$ As Gühring et al., (2020, Proof of Lemma C.4, Step 2) showed, the conditions of the Bramble-Hilbert-Lemma (Brenner & Scott,, 2008, Theorem 4.3.8) are satisfied. Hence for $k\in\{0,1\}$
$|\tilde{f}-c_{m,f}|_{W^{k,\infty}(\Omega_{m,N})}\leq C_{1}\Big{(}\frac{2\sqrt{d}}{N}\Big{)}^{1-k}|\tilde{f}|_{W^{1,\infty}(\Omega_{m;n})}\leq C_{2}\Big{(}\frac{1}{N}\Big{)}^{1-k}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m;n})}.$
Now using $\phi_{m}$ as defined above, we get
$\left\|\phi_{m}\Big{(}\tilde{f}-c_{f,m}\Big{)}\right\|_{\infty,\Omega_{m,N}}$ $\leq\left\|\phi_{m}\right\|_{\infty,\Omega_{m,N}}\cdot\left\|\tilde{f}-c_{f,m}\right\|_{\infty,\Omega_{m,N}}\leq C_{2}\frac{1}{N}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}$ (24)
Due to the product inequality for weak derivatives (Gühring et al.,, 2020, Lemma B.6) there is a constant $C^{\prime}>0$ such that the supremum norm of the weak derivative is bounded by
$\left|\phi_{m}\Big{(}\tilde{f}-c_{f,m}\Big{)}\right|_{W^{1,\infty}(\Omega_{m,N})}\leq$ $C^{\prime}|\phi_{m}|_{W^{1,\infty}(\Omega_{m,N})}\cdot\left\|\tilde{f}-c_{f,m}\right\|_{\infty,\Omega_{m,N}}$
$+C^{\prime}\left\|\phi_{m}\right\|_{\infty,\Omega_{m,N}}\cdot|\tilde{f}-c_{f,m}|_{W^{1,\infty}(\Omega_{m,N})}$
$\leq$ $C^{\prime}\cdot cN\cdot C_{2}\frac{1}{N}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}+C^{\prime}\cdot C_{3}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}$
$=$ $C_{4}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}.$ (25)
Combining (24) and (25) we get
$\left\|\phi_{m}\Big{(}\tilde{f}-c_{f,m}\Big{)}\right\|_{W^{1,\infty}(\Omega_{m,N})}\leq C_{5}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}.$
Step 3 (Global estimate in $\|\cdot\|_{W^{k,p}}$): As $\sum_{m\in\{0,...,N\}^{d}}\phi_{m}=1,$ we have that
$\tilde{f}(x)=\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}(x)\tilde{f}(x),\quad\text{ for a.e. }x\in(0,1)^{d}.$
As $\tilde{f}\big{|}_{(0,1)^{d}}=f$ we have for $k\in\{0,1\}$
$\Big{\|}f-\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}c_{f,m}\Big{\|}_{W^{k,\infty}((0,1)^{d})}$ $=\Big{\|}\tilde{f}-\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}c_{f,m}\Big{\|}_{W^{k,\infty}((0,1)^{d})}$
$=\Big{\|}\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}\big{(}\tilde{f}-c_{f,m}\big{)}\Big{\|}_{W^{k,\infty}((0,1)^{d})}$
$\leq\sup_{\widetilde{m}\in\{0,\ldots,N\}^{d}}\ \Big{\|}\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}\big{(}\tilde{f}-c_{f,m}\big{)}\Big{\|}_{W^{k,\infty}(\Omega_{\widetilde{m},N})}$ (26)
where the last step follows from $(0,1)^{d}\subset\bigcup_{\widetilde{m}\in\{0,\ldots,N\}^{d}}\Omega_{\widetilde{m},N}$. Now we obtain for each $\widetilde{m}\in\{0,\ldots,N\}^{d}$ using (23), (24) and (25)
$\Big{\|}\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}\big{(}\tilde{f}-c_{f,m}\big{)}\Big{\|}_{W^{k,\infty}(\Omega_{\tilde{m},N})}$ $\leq\sup_{\begin{subarray}{c}m\in\{0,\ldots,N\}^{d}\\ |m-\widetilde{m}|_{\infty}\leq 1\end{subarray}}\big{\|}\phi_{m}\big{(}\tilde{f}-c_{f,m}\big{)}\big{\|}_{W^{k,\infty}(\Omega_{\tilde{m},N})}$
$\leq\sup_{\begin{subarray}{c}m\in\{0,\ldots,N\}^{d}\\ |m-\widetilde{m}|_{\infty}\leq 1\end{subarray}}\big{\|}\phi_{m}\big{(}\tilde{f}-c_{f,m}\big{)}\big{\|}_{W^{k,\infty}(\Omega_{m,N})}$
$\leq C_{6}\Big{(}\frac{1}{N}\Big{)}^{1-k}\sup_{\begin{subarray}{c}m\in\{0,\ldots,N\}^{d},\\ |m-\widetilde{m}|_{\infty}\leq 1\end{subarray}}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}.$
Plugging this into (26), we obtain for $k\in\{0,1\}$
$\Big{\|}f-\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}c_{f,m}\Big{\|}_{W^{k,\infty}((0,1)^{d})}$ $\leq C_{6}\Big{(}\frac{1}{N}\Big{)}^{(1-k)}\sup_{\tilde{m}\in\{0,\ldots,N\}^{d}}\Big{(}\sup_{\begin{subarray}{c}m\in\{0,\ldots,N\}^{d},\\ |m-\widetilde{m}|_{\infty}\leq 1\end{subarray}}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}\Big{)}$ (29)
$\leq C_{7}\Big{(}\frac{1}{N}\Big{)}^{(1-k)}\sup_{\widetilde{m}\in\{0,\ldots,N\}^{d}}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{\widetilde{m},N})}$
$\leq C_{8}\Big{(}\frac{1}{N}\Big{)}^{(1-k)}\|\tilde{f}\|_{W^{1,\infty}(\mathbb{R}^{d})}$
$\leq C_{9}\Big{(}\frac{1}{N}\Big{)}^{(1-k)}\|f\|_{W^{1,\infty}((0,1)^{d})}.$ (30)
Step 4 (Interpolation): Define the linear operators $T_{0}\colon W^{1,\infty}((0,1)^{d})\rightarrow L^{\infty}((0,1)^{d}),T_{\alpha}\colon W^{1,\infty}((0,1)^{d})\rightarrow\mathcal{H}^{\alpha}((0,1)^{d})$ and $T_{1}\colon W^{1,\infty}((0,1)^{d})\rightarrow W^{1,\infty}((0,1)^{d})$ via
$T_{k}(f)=f-\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}c_{f,m},\quad k\in\{0,\alpha,1\}$
Note that the linearity follows from the definition of the constants $c_{f,m}.$ Using Lunardi, (2018, Theorem 1.6), for the nontrivial interpolation couple see (Lunardi,, 2018, p.11 f.), leads to
$\|T_{\alpha}\|\leq\|T_{0}\|^{1-\alpha}\|T_{1}\|^{\alpha}.$
Note that $\|\cdot\|_{\mathcal{H}^{\alpha}}$ is equivalent to $\|\cdot\|_{W^{s,\infty}(\Omega)}$ in Gühring et al., (2020). Using (30) we conclude
$\Big{\|}f-\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}c_{f,m}\Big{\|}_{\mathcal{H}^{\alpha}}$ $\leq C_{10}\Big{(}\frac{1}{N}\Big{)}^{1-\alpha}\|f\|_{W^{1,\infty}}.\qed$
Now we want to approximate the function $\sum_{m\in\{0,...,N\}^{d}}c_{f,m}\phi_{m}$ in Hölder norm using a ReLU network.
Lemma 7.4.
For any $\varepsilon\in(0,1/2)$ there is a neural network $\Phi_{\varepsilon}$ with ReLU activation function such that for $(c_{f,m})_{m}$ from Lemma 7.3, there is a constant $C>0$ such that
$\Big{\|}\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}c_{f,m}-\Phi_{\varepsilon}\Big{\|}_{\mathcal{H}^{\alpha}}\leq C\|f\|_{W^{1,\infty}}N^{\alpha}\varepsilon,$
the number of layers is at most $\lceil C_{1}\log_{2}(\varepsilon^{-1})\rceil,$ the number of nonzero weights is at most $\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\varepsilon^{-1})\rceil$ and the number of neurons is at most $\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\varepsilon^{-1})\vee\log_{2}(\varepsilon^{-1}))\rceil,$ with $C_{1},C_{2}$ and $C_{3}$ from Lemma 7.2.
Proof.
From Lemma 7.2 we know that there are neural networks $\Psi_{\varepsilon,m}$ with at most $\lceil C_{1}\log_{2}(\varepsilon^{-1})\rceil$ layers, $\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\varepsilon^{-1})\rceil$ nonzero weights and $\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\varepsilon^{-1})\vee\log_{2}(\varepsilon^{-1}))\rceil$ neurons that approximate $\phi_{m}$ such that for $k\in\{0,1\}$
$\left\|\phi_{m}-\Psi_{\varepsilon,m}\right\|_{W^{k,\infty}}\leq c^{\prime}N^{k}\varepsilon.$
Now we parallelize these networks and multiply with the coefficients $c_{f,m}$ afterwards. Hereby, we construct a network $\Phi_{\varepsilon}$ with $1+\lceil C_{1}\log_{2}(\varepsilon^{-1})\rceil$ layers, $N^{d}+\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\varepsilon^{-1})\rceil$ nonzero weights and $1+\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\varepsilon^{-1})\vee\log_{2}(\varepsilon^{-1}))\rceil$ neurons such that
$\Phi_{\varepsilon}=\sum_{m\in\{0,...,N\}^{d}}c_{f,m}\Psi_{\varepsilon,m}.$ (31)
For each $m\in\{0,...,N\}^{d}$ denote $\Omega_{m,N}=B_{\frac{1}{N},|\cdot|_{\infty}}(\frac{m}{N})$ as above. For $k\in\{0,1\}$ we get
$\Big{\|}\Phi_{\varepsilon}-\sum_{m\in\{0,\ldots,N\}^{d}}c_{f,m}\phi_{m}\Big{\|}_{W^{k,\infty}((0,1)^{d})}=\Big{\|}\sum_{m\in\{0,\ldots,N\}^{d}}c_{f,m}(\Psi_{\varepsilon,m}-\phi_{m})\Big{\|}_{W^{k,\infty}((0,1)^{d})}$
$\qquad\qquad\leq\sup_{\widetilde{m}\in\{0,\ldots,N\}^{d}}\Big{\|}\sum_{m\in\{0,\ldots,N\}^{d}}c_{f,m}(\Psi_{\varepsilon,m}-\phi_{m})\Big{\|}_{W^{k,\infty}(\Omega_{\widetilde{m},N}\cap(0,1)^{d})}$
$\qquad\qquad\leq 3^{d}\sup_{\widetilde{m}\in\{0,\ldots,N\}^{d}}\sup_{m\in\{0,\ldots,N\}^{d}}\Big{\|}c_{f,m}(\Psi_{\varepsilon,m}-\phi_{m})\Big{\|}_{W^{k,\infty}(\Omega_{\widetilde{m},N}\cap(0,1)^{d})}$
$\qquad\qquad\leq 3^{d}\sup_{\widetilde{m}\in\{0,\ldots,N\}^{d}}\sup_{m\in\{0,\ldots,N\}^{d}}|c_{f,m}|\Big{\|}(\Psi_{\varepsilon,m}-\phi_{m})\Big{\|}_{W^{k,\infty}(\Omega_{\widetilde{m},N}\cap(0,1)^{d})}$
$\qquad\qquad\leq 3^{d}\sup_{\widetilde{m}\in\{0,\ldots,N\}^{d}}\sup_{m\in\{0,\ldots,N\}^{d}}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}\|\Psi_{\varepsilon,m}-\phi_{m}\|_{W^{k,\infty}(\Omega_{\widetilde{m},N}\cap(0,1)^{d})}$
$\qquad\qquad\leq CN^{k}\varepsilon\|f\|_{W^{1,\infty}}.$
The second to last inequality follows from the fact that on $\Omega_{\widetilde{m},N}$ is within the support of $\phi_{m}$ only for $|m-\widetilde{m}|_{\infty}\leq 1.$ The last inequality follows from (31) and the continuity of the extension operator, see (Stein,, 1970, Theorem 5). As in Step 4 of Lemma 7.3, we conclude using Lunardi, (2018, Theorem 1.6)
$\Big{\|}\Phi_{\varepsilon}-\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}c_{f,m}\Big{\|}_{\mathcal{H}^{\alpha}}\leq CN^{\alpha}\varepsilon\|f\|_{W^{1,\infty}}.\qed$
Now we are ready to proof Theorem 5.3.
Proof of Theorem 5.3.
Combining Lemma 7.3 and Lemma 7.4 with $\|f\|_{W^{1,\infty}}\leq B$ yields for a constant $C>0$ for any $\tilde{\varepsilon}\in(0,1/2)$ that
$\|f-\Phi_{\tilde{\varepsilon}}\|_{\mathcal{H}^{\alpha}}$ $\leq\Big{\|}f-\sum_{m\in\{0,\ldots,N\}^{d}}c_{f,m}\phi_{m}\Big{\|}_{\mathcal{H}^{\alpha}}+\Big{\|}\sum_{m\in\{0,\ldots,N\}^{d}}c_{f,m}\phi_{m}-\Phi_{\tilde{\varepsilon}}\Big{\|}_{\mathcal{H}^{\alpha}}$
$\leq CB\Big{(}\Big{(}\frac{1}{N}\Big{)}^{1-\alpha}+N^{\alpha}\tilde{\varepsilon}\Big{)},$ (32)
where $\tilde{\varepsilon}$ determines the approximation accuracy in Lemma 7.4. For
$N:=\left\lceil\Big{(}\frac{\varepsilon}{2CB}\Big{)}^{-1/(1-\alpha)}\right\rceil,$
we get for the first term in (32)
$\Big{(}\frac{1}{N}\Big{)}^{1-\alpha}\leq\frac{\varepsilon}{2CB}.$
Choosing
$\tilde{\varepsilon}=\frac{\varepsilon}{2CB}\Big{(}\Big{(}\frac{\varepsilon}{2CB}\Big{)}^{-\frac{1}{1-\alpha}}+1\Big{)}^{-\alpha}$ (33)
leads to
$\|f-\Phi_{\tilde{\varepsilon}}\|_{\mathcal{H}^{\alpha}}$ $\leq\varepsilon.$
From Lemma 7.2 we know that there is a ReLU network with no more than $1+\lceil C_{1}\log_{2}(\tilde{\varepsilon}^{-1})\rceil$ layers, $N^{d}+\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\tilde{\varepsilon}^{-1})\rceil$ nonzero weights and $1+\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\tilde{\varepsilon}^{-1})\vee\log_{2}(\tilde{\varepsilon}^{-1}))\rceil$ neurons with the required properties. Inserting (33) and assuming $CB>\frac{1}{2}$ yields
$\log_{2}(\tilde{\varepsilon}^{-1})$ $\leq\log_{2}\Big{(}\frac{2CB}{\varepsilon}2^{\alpha}\Big{(}\frac{\varepsilon}{2CB}\Big{)}^{-\frac{\alpha}{1-\alpha}}\Big{)}\leq C^{\prime}\log_{2}(\varepsilon^{-\frac{1}{1-\alpha}}).$
Thus there are $C^{\prime},C^{\prime\prime}$ and $C^{\prime\prime\prime}$ such that the ReLU network has no more than $1+\lceil C^{\prime}\log_{2}(\varepsilon^{-\frac{1}{1-\alpha}})\rceil$ layers, $\lceil C^{\prime\prime}\varepsilon^{-\frac{d}{1-\alpha}}\log^{2}_{2}(\varepsilon^{-\frac{1}{1-\alpha}})\rceil$ nonzero weights and $1+\lceil C^{\prime\prime\prime}\varepsilon^{-\frac{d}{1-\alpha}}(\log_{2}^{2}(\varepsilon^{-\frac{1}{1-\alpha}})\vee\log_{2}(\varepsilon^{-\frac{1}{1-\alpha}}))\rceil$ neurons.
Since $f\in\operatorname*{Lip}(L,B)\subseteq\mathcal{H}^{\alpha}(\Gamma)$ for $\Gamma=\max(L,2B)$, we conclude
$\|\Phi_{\tilde{\varepsilon}}\|_{\mathcal{H}^{\alpha}}\leq\|f\|_{\mathcal{H}^{\alpha}}+\|\Phi_{\tilde{\varepsilon}}-f\|_{\mathcal{H}^{\alpha}}\leq\Gamma+\varepsilon.$
Corollary 5.4 is a straightforward combination of Theorem 5.2 and Theorem 5.3. ∎
7.6 Proof of Theorem 6.1
Proof of Theorem 6.1.
We get with the same reasoning as in the proof of (32)
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})$ $\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\mathsf{W}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})+2\inf_{W^{\prime}\in\mathcal{W}}\sup_{W\in\operatorname{Lip}(1,B)}\|W-W^{\prime}\|_{\infty}$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\mathsf{W}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})+2\inf_{W^{\prime}\in\mathcal{W}}\sup_{W\in\operatorname{Lip}(1,B)}\|W-W^{\prime}\|_{\infty}$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\mathsf{W}_{\mathcal{H}^{\alpha}(\Gamma)}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})+2\inf_{W^{\prime}\in\mathcal{W}}\sup_{W\in\operatorname{Lip}(1,B)}\|W-W^{\prime}\|_{\infty}.$
The bound on $\mathsf{W}_{\mathcal{H}^{\alpha}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})$ depending on the intrinsic dimension $d^{*}$ was already derived in Theorem 5.2. Along Corollary 4.2 we obtain the rate depending on the ambient dimension $d.$
Note that the Wasserstein distance is invariant to additive shifts, hence we can add the assumption $W(0)=0$ without changing the optimal value. As $\mathcal{X}$ is bounded, there exists a $B>0$ depending on the norm used on $\mathcal{X}.$ In any case $B=d$ is sufficient. To apply Theorem 5.3 we need the Hölder constant of the network to be greater or equal than $2B.$
Appendix A Calculations for Example 3.3
For the Wasserstein distance we get $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})=\gamma.$ The Vanilla GAN distance using all Lipschitz $L$ affine functions as discriminator yields in this example $\mathsf{V}_{a\cdot+b}(\mathbb{P},\mathbb{Q})=\max_{\begin{subarray}{c}a,b\in\mathbb{R},\\ |a|\leq L\end{subarray}}f(a,b)$ for
$f(a,b)\coloneqq\frac{1}{2}\big{(}-\log\big{(}1+e^{-a\gamma-b}\big{)}-\log\big{(}1+e^{-a(\gamma+\varepsilon)-b}\big{)}-\log\big{(}1+e^{b}\big{)}-\log\big{(}1+e^{a\varepsilon+b}\big{)}\big{)}+\log(4).$
Standard calculus yields for fixed $a$ the unique maximizer $b^{*}=-\frac{a(\varepsilon+\gamma)}{2}$ and
$f(a,b^{*})=-\log\Big{(}1+e^{-\frac{a(\varepsilon+\gamma)}{2}}\Big{)}-\log\Big{(}1+e^{\frac{a(\varepsilon-\gamma)}{2}}\Big{)}+\log(4).$
Since
$\frac{\partial}{\partial a}f(a,b^{*})=\frac{\varepsilon+\gamma}{2(e^{\frac{a(\gamma+\varepsilon)}{2}}+1)}-\frac{\varepsilon-\gamma}{2(e^{-\frac{a(\varepsilon-\gamma)}{2}}+1)},$
for $\varepsilon\leq\gamma,$ the maximizing $a$ is maximal $a^{*}=L.$ This coincides with the intuitive choice: as the support of $\mathbb{P}^{X}$ and the support of $\mathbb{P}^{Y}$ can be separated by a single point on $\mathbb{R},$ we expect the optimal discriminator to be affine linear. Standard calculus yields the linear upper and lower bound for $\varepsilon=\frac{1}{4}$.
For $\varepsilon>\gamma,$ the unrestricted maximizing $a^{*}$ solves the equation
$(\varepsilon-\gamma)e^{\frac{a^{*}(\varepsilon+\gamma)}{2}}-(\varepsilon+\gamma)e^{-\frac{a^{*}(\varepsilon-\gamma)}{2}}=2\gamma.$
While there is no closed form solution, a numerical approximation (for $\varepsilon=\frac{1}{4}$) yields for $\gamma<\varepsilon$ and $L>16$ such that $a^{*}$ is feasible
$\frac{\mathsf{W}_{1}(\mathbb{P}^{X},\mathbb{P}^{Y})^{2}}{2}\leq\mathsf{V}_{a\cdot+b}(\mathbb{P}^{X},\mathbb{P}^{Y})\leq a\cdot\mathsf{W}_{1}(\mathbb{P}^{X},\mathbb{P}^{Y})^{2}.$